### v2

# Learning to Control Unknown Strongly Monotone Games

Abstract
Consider a game where the players’ utility functions include a reward function and a linear term for each dimension, with coefficients that are controlled by the manager. We assume that the game is strongly monotone, so gradient play converges to a unique Nash equilibrium (NE). The NE is typically globally inefficient. The global performance at NE can be improved by imposing linear constraints on the NE. We therefore want the manager to pick the controlled coefficients that impose the desired constraint on the NE. However, this requires knowing the players’ reward functions and action sets. Obtaining this game information is infeasible in a large-scale network and violates user privacy. To overcome this, we propose a simple algorithm that learns to shift the NE to meet the linear constraints by adjusting the controlled coefficients online. Our algorithm only requires the linear constraints violation as feedback and does not need to know the reward functions or the action sets. We prove that our algorithm converges with probability 1 to the set of NE that satisfy target linear constraints. We then prove an L2 convergence rate of near-$O(t^{-1/4})$.
1 Introduction
Large-scale networks such as in communication, transportation, energy, and computing consist of myriad of agents taking local decisions. Agents do not see the bigger picture therefore optimize their local objectives instead. The result is often an inefficient equilibrium. Hypothetically, a centralized authority that knows all network parameters can instruct the agents how to act to optimize a global objective. However, such a centralized architecture is infeasible for large-scale networks and also violates the users’ privacy. As a result, there has been a surge of interest in distributed multiagent algorithms [1, 2, 3, 4, 5].
Game theory is a natural way to tackle distributed decision-making [6, 7]. The game structure formalizes the effect of the interaction between users. The players may be cooperative, or they may be selfishly interested in maximizing their reward. In large-scale systems, obtaining performance guarantees is challenging even with cooperative players since they do not know the reward functions of their peers. A Nash Equilibrium (NE) aims to predict the outcome of such a repeated interaction. In monotone games [8, 9, 10], which are our focus here, simple distributed algorithms such as gradient descent of players on their reward functions are guaranteed to converge to a NE.
A NE does not optimize global objectives and can lead to poor performance. By controlling parameters in the utility functions of the players, the manager can aim to change the game such that the NE is efficient. In general, setting utility parameters to optimize performance at the NE is intractable. Therefore, in this paper, we focus on the task of imposing linear constraints on the resulting NE of the game the manager controls. Linear constraint are common in many optimization formulations. When optimizing a global quadratic cost, linear constraint can describe the first-order optimality condition for the NE. In resource allocation, linear constraints can describe what should be the ideal loads at each resource at NE.
Even when setting utility parameters to impose linear constraints at the NE is tractable, it requires knowing the reward functions of players and their action sets. This would require the manager to collect this information from the players, and then solve a large-scale optimization problem. This centralized approach violates the privacy of the users and is infeasible in large-scale networks. Even worse, the manager would need to collect the parameters periodically if they are time-varying, and then solve the large-scale optimization problem again.
To converge to a NE that meets target linear constraints in an unknown strongly monotone game, we propose an online learning approach. In our algorithm, the manager adjusts the controlled coefficients online and receives the constraint violation as feedback. Such feedback is easy to monitor in practice and maintains user privacy. Our algorithm is based on two time-scale stochastic approximation, where the agents update their actions in the faster time-scale and the manager updates the controlled coefficients in the slower time-scale.
We prove that our algorithm converges to the set of NE that satisfy the target linear constraints with probability 1. We further show that our algorithm has a mean square error rate of near-$O(t^{-1/4})$ for the linear constraint. Due to the convex and compact action sets, our slower time-scale is a fixed point iteration with a non-expansive mapping instead of the stronger contractive mapping typically assumed to analyze convergence rates [11, 12]. Apart from stochastic approximation, we employ tools from variational inequalities and the Krasnosel’skiĭ–Mann (KM) algorithm in our analysis. Therefore, we provide a theoretically principled game control scheme for strongly monotone games.
1.1 Related Work
Distributed optimization [13, 14] is concerned with optimizing a target objective distributedly over a network of agents. However, these agents are non-interacting computational units so there is no game structure between them. To cast the game as a distributed optimization with the action profile as the variable, each agent would need to know the effect of other agents on its reward function, which is infeasible. Moreover, distributed optimization entails significant communication overhead between agents.
Distributed algorithms that converge to NE have been studied in [15, 16, 17], including in monotone games [18, 19, 20]. The players in our game use simple stochastic gradient play to converge to NE. However, if the game is uncontrolled, the NE is typically globally inefficient. Using our game control scheme, the manager learns to steer the game into an efficient NE (i.e., a NE that satisfies target linear constraints).
Compared to mechanism design approaches [21, 22], we focus on cooperative or myopic players and our manager does not elicit private information from the players but only observes the constraint violation.
Our approach resembles a Stackelberg game [23, 24, 25], where the manager is the leader and the players are the followers (usually a single follower is considered). However, game-theoretic algorithms converge to a NE that can be globally inefficient. Our work provides a mechanism that provably shifts the unique NE to a point that satisfies target linear constraints that can force the NE to be efficient (e.g., first-order optimality condition for a global quadratic objective).
Our learning framework can be viewed as a bandit learning problem where the manager is playing against a bandit that is not stochastic nor fully adversarial [26, 27]. Instead, the rewards of this bandit are generated by the game dynamics. This “game bandit” introduces a new type of noise. The goal of the manager is to steer the unique NE to a point that satisifes target linear constraints. Hence, the type of feedback the manager needs concerns the behavior at NE. However, the dynamics are not at NE every turn. In fact, by changing the prices the manager perturbs the system which complicates and delays the convergence to the unique NE. Therefore, the main noise in this “game bandit” is an equilibrium noise that stems from the distance of the dynamics from the time-varying NE.
Our problem can be viewed as a Generalized Nash Equilibrium Problem (GNEP), where players seek to converge to a NE subject to coupled constraints [28]. Efficient computation of GNE in strongly monotone games in the noiseless setting has been widely studied [29, 30, 31]. Stochastic convergence to GNEs has recently been studied under different assumptions. In [32] players share their actions with others, and in [33] the reward function is assumed to be linear. In [34], stochastic convergence to GNE is studied more generally, resulting in a centralized algorithm where the weighted empirical average of the actions is shown to converge in expectation. We provide a distributed stochastic learning scheme that converges in probability 1 to GNE with linear coupled constraints in strongly monotone games without communication between players.
Our work belongs to the literature on intervention and control of games [35, 36, 37, 38, 39, 40, 41, 42, 43]. In [44] it was assumed that the manager knows the reward functions of the players, and can compute the desired NE. This is also the assumption in [45, 39], that considered the general approach of game-theoretic control. In [35], an aggregative game where the cost is a linear function of the aggregated action was considered. It was assumed that the manager can control the total loads the players observe, which can deviate from the true total loads. In [46, 47, 48], a similar approach to ours was considered for discrete congestion games, where the reward function of all players is the same convex function of the number of players that share the chosen resource. The congestion game considered there is a special case of monotone games that allow us to model far more general interactions and action sets. For example, it allows for models when not only the number of players that share a resource matters but also their identity (e.g., geometric locations). Additionally, the pricing scheme in [46, 47, 48] requires the manager to know the reward functions. Control of dynamic Markov potential games was considered in [40], where it was assumed that the manager knows the players’ reward functions and can simulate the game.
Some works obtained convergence rates for contractive two time-scale stochastic approximation [49, 11]. These works assume that both the time-scales are fixed point iterations with contractive mappings. We consider a more general case where the slower timescale has a non-expansive mapping.
An earlier conference paper [50] was limited to load balancing in resource allocation. This paper’s framework is similar to that in [50], but is extended to linear constraint in strongly monotone games. The NE guarantees here are of the form $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$ instead of just $\sum_{n=1}^{N}\boldsymbol{x}_{n}\left(\boldsymbol{\alpha}^{*}\right)=\boldsymbol{\ell^{*}}$. Nevertheless, the results here are completely novel and independent of [50]. Our convergence guarantees are with probability 1 instead of in L2 like in [50]. Unlike [50], we provide a finite time convergence error bound. Our novel analysis is based on two time-scale stochastic approximation with non-expansive mappings.
1.2 Outline and Notation
The rest of the paper is organized as follows. Section 2 formulates our game control problem and details our assumptions. Subsection 2.1 details two applications that fall under our general model. Section 3 describes the game control algorithm and states our main results regarding its performance guarantees. Section 4 breaks the proofs of our main results into lemmas. Section 5 provides numerical simulations for the application described earlier. Finally, Section 6 concludes the paper and proposes future research directions.
[…]
2 Problem Formulation
Consider a set of players $\mathcal{N}=\{1,2,\ldots,N\}$, where player $n$ chooses an action $\boldsymbol{x}_{n}=(x_{n}^{1},\ldots,x_{n}^{d})\in\mathcal{X}_{n}$. Here $\mathcal{X}_{n}\subset\mathbb{R}^{d}_{+}$ is a convex and compact set such that $\boldsymbol{0}\in\mathcal{X}_{n}$. Let $\mathcal{X}=\mathcal{X}_{1}\times\ldots\times\mathcal{X}_{n}$. We use $\boldsymbol{x}\in\mathcal{X}$ to denote the concatenation of all action vectors. The utility for each player is defined as:
$u_{n}(\boldsymbol{x};\boldsymbol{\beta})=r_{n}(\boldsymbol{x})-\sum_{i=1}^{d}\beta_{n}^{i}x_{n}^{i}.$ (1)
Here $r_{n}(\boldsymbol{x})$ is the reward function for player $n$. We assume that $r_{n}(\boldsymbol{x})$ is twice continuously differentiable for each $n$ and is greater than or equal to 00 for all $\boldsymbol{x}\in\mathcal{X}$. The vector $\boldsymbol{\beta}\in\mathbb{R}^{Nd}$ denotes the ‘controlled coefficients’ controlled by the game manager. We use the same vector notation for $\boldsymbol{\beta}$ and $\boldsymbol{\beta}_{n}$.
Figure 1: Control of an unknown game
The key assumption about the game structure is that it is a strongly monotone game. For this, we define the gradient operator $F(\cdot):\mathbb{R}^{Nd}\mapsto\mathbb{R}^{Nd}$ as
$F(\boldsymbol{x})\coloneqq(\nabla_{\boldsymbol{x}_{1}}r_{1}(\boldsymbol{x}_{1},\boldsymbol{x}_{-1}),\ldots,\nabla_{\boldsymbol{x}_{N}}r_{N}(\boldsymbol{x}_{N},\boldsymbol{x}_{-N})).$ (2)
Assumption 1.
There exists $\mu>0$ such that for all $\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
$\langle\boldsymbol{y}-\boldsymbol{x},F(\boldsymbol{y})-F(\boldsymbol{x})\rangle\leq-\mu\|\boldsymbol{y}-\boldsymbol{x}\|^{2}.$ (3)
The above assumption means that the game is a strongly monotone game with parameter $\mu>0$. Monotone games are a well-studied class of games [51]. The linear term in (1) does not affect the strongly monotone property of the game since the gradient vector becomes $F(\boldsymbol{x})-\boldsymbol{\beta}$. Since it is a strongly monotone game, there exists a unique pure NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$ corresponding to each vector $\boldsymbol{\beta}$, defined as:
Definition 1.
A strategy profile $(\boldsymbol{x^{*}}_{n}(\boldsymbol{\beta}),\boldsymbol{x^{*}}_{-n}(\boldsymbol{\beta}))$ is a pure Nash equilibrium (NE) corresponding to coefficients $\boldsymbol{\beta}$, if $u_{n}(\boldsymbol{x}_{n}^{\boldsymbol{*}}(\boldsymbol{\beta}),\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})\geq u_{n}(\boldsymbol{x}_{n},\boldsymbol{x}_{-n}^{\boldsymbol{*}}(\boldsymbol{\beta});\boldsymbol{\beta})$, for all $\boldsymbol{x}_{n}\in\mathcal{X}$ and all $n\in\mathcal{N}$.
We make the following additional assumption about the game, which says that the gradient function is Lipschitz continuous. It is merely a technical assumption, commonly taken along with the strongly monotone condition [52, 53].
Assumption 2.
There exists $L>0$ such that for all $\boldsymbol{x},\boldsymbol{y}\in\mathcal{X}$,
$\|F(\boldsymbol{y})-F(\boldsymbol{x})\|\leq L\|\boldsymbol{y}-\boldsymbol{x}\|.$ (4)
Our game evolves in discrete time, so at time $t$ players take actions $\boldsymbol{x}_{n,t}$ given the controlled coefficients $\boldsymbol{\beta}_{n,t}$. The players update their actions using gradient ascent and the controlled coefficients are adjusted based on the manager’s objective (discussed below (6)). Since the game is strongly monotone, if the controlled coefficients are fixed at $\boldsymbol{\beta}$ (i.e., uncontrolled game), the players will converge to the NE $\boldsymbol{x^{*}}(\boldsymbol{\beta})$.
In many scenarios, the players do not fully know their reward functions but instead learn them on the fly based on the reward values they receive (i.e., bandit feedback) or other data they collect. The gradient can then be approximated from this data using regression or other supervised machine learning techniques [54]. These learning schemes make the players’ actions stochastic which adds another type of noise to our analysis. We therefore assume that agents observe noisy variants of their gradients, where the noise is a martingale difference sequence with a bounded second moment.
Assumption 3.
Each player $n$ observes $\boldsymbol{g}_{n,t-1}$ at time $t$, where $\boldsymbol{g}_{n,t-1}=\nabla_{\boldsymbol{x}_{n,t-1}}r_{n}(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{n,t}$. Then $\boldsymbol{g}_{t-1}=F(\boldsymbol{x}_{t-1})+\boldsymbol{M}_{t}$, where $\boldsymbol{M}_{t}$ satisfies
$\mathbb{E}[\boldsymbol{M}_{t}\mid\mathcal{F}_{t-1}]=0\;\;\;\textrm{and}\;\;\;\mathbb{E}[\|\boldsymbol{M}_{t}\|^{2}\mid\mathcal{F}_{t-1}]\leq M_{c},$ (5)
for some $M_{c}>0$ and a filtration $\mathcal{F}_{t}$ that summarizes the past:
$\mathcal{F}_{t}=\sigma(\boldsymbol{x}_{\tau},\boldsymbol{\beta}_{\tau}\mid\forall\tau\leq t).$
The objective of the manager is to steer the players’ NE towards a point that satisfies a set of $K$ linear constraints, which can together be represented in vector form as
$\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}.$ (6)
Here $\boldsymbol{A}\in\mathbb{R}^{K\times Nd}$ and $\boldsymbol{\ell^{*}}\in\mathbb{R}^{K}$. We assume that, at each time $t$, the manager only observes the constraint violation, i.e., the vector $(\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}})$. These linear constraints can have different meanings for different applications. For resource allocation, the players’ actions are how much to use of each resource and the manager wants to meet the target loads. When the global objective is a quadratic cost, the constraint violation vector is just the gradient of the global quadratic cost.
2.1 Applications
In this section, we show how our framework can be applied in two different applications.
2.1.1 Quadratic Global Objective
For a positive semi-definite matrix $\boldsymbol{G}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{\rho}\in\mathbb{R}^{Nd}$, the controller can set a global cost of the form $\Phi\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$. All the action profiles that are local optima of $\Phi\left(\boldsymbol{x}\right)$ satisfy $\nabla\Phi\left(\boldsymbol{x}\right)=2\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}=0$. Hence, using our approach, we can set $\boldsymbol{A}=2\boldsymbol{G}$ and $\boldsymbol{\ell^{*}}=-\boldsymbol{\rho}$ to reach a global optimum.
2.1.2 Weighted Resource Allocation Games
In resource allocation, high congestion on a single resource increases the operation costs for the manager, can lead to system failures, and incurs suboptimal performance for the users. Therefore, load balancing is a key control objective for the manager that supervises the resource allocation protocol. In the electricity grid, the manager can be the utility company running Demand Side Management (DSM) to cut the production costs of energy at peak hours. In wireless networks, the manager can be an access point coordinating a protocol that minimizes the power consumption of the devices and the interference in the network. Other examples are load balancing in data-centers [55] and control of parking resources [56]. For distributed resource allocation, designing an efficient load balancing protocol becomes far more challenging.
In weighted resource allocation, each dimension of a player’s action represents how much they use from each of the $d$ resources. We have $d$ linear constraints i.e., $K=d$ and each represents a target load for each resource $i$, i.e.,
[…]
The manager can set the target loads $\boldsymbol{l}^{*}$ to optimize application-specific objectives such as system efficiency, regulations, and operation costs. The controlled coefficients $\boldsymbol{\beta}$ represent incentives to use a resource. The weights in $\boldsymbol{A}$ are based on how each agent’s use affects the resource.
A special case is resource allocation without weights, studied in [50]. Generalizing to a weighted sum of actions allows modeling of power control in wireless networks and load balancing in data centers. In wireless networks, the interference that the manager measures is a weighted sum of the transmission powers of the players, where the weights are the channel gains. In data centers, different users send jobs of different sizes to different servers (resources), so the load is a weighted sum of their actions (how many jobs to send where).
3 Game Control Algorithm
Algorithm 1 Online Game Control with Bandit Feedback
Initialization: Let $\boldsymbol{x}_{0}\in\mathcal{X}$ and $\boldsymbol{\alpha}_{0}\in\mathbb{R}^{K}$.
Input: Target vector $\boldsymbol{\ell^{*}}$.
For each turn $t\geq 1$ do
  1. 1.
    The manager observes the vector $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$.
  2. 2.
    Each player $n$ computes $\boldsymbol{\beta}_{n,t-1}=\boldsymbol{A}^{\intercal}_{n}\boldsymbol{\alpha}_{t-1}$ and updates its action using gradient ascent:
    $\boldsymbol{x}_{n,t}=\Pi_{\mathcal{X}_{n}}\left(\boldsymbol{x}_{n,t-1}+\eta_{t-1}\left(\boldsymbol{g}_{n,t-1}-\boldsymbol{\beta}_{n,t-1}\right)\right)$ (7)
    where $\Pi_{\mathcal{X}_{n}}$ is the Euclidean projection into $\mathcal{X}_{n}$.
  3. 3.
    The manager updates the controlled input using
    $\boldsymbol{\alpha}_{t}=\boldsymbol{\alpha}_{t-1}+\epsilon_{t-1}(\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}).$ (8)
  4. 4.
    The manager broadcasts $\boldsymbol{\alpha}_{t}$ to the agents.
End
Algorithm 1 details the combined scheme of the manager’s iteration and the players’ behavior (modeled or designed). We discuss the key details of the algorithm in this section.
The manager controls the controlled coefficients $\boldsymbol{\beta}$ in order to steer the dynamics towards a desirable point. But instead of directly performing computations on $\boldsymbol{\beta}\in\mathbb{R}^{Nd}$ and communicating it to agents, the manager uses the “control input” $\boldsymbol{\alpha}\in\mathbb{R}^{K}$, where $K$ is the number of linear constraints, which can be much smaller than $Nd$ depending on the application. After computing $\boldsymbol{\alpha}_{t}$ at time $t$, the manager broadcasts $\boldsymbol{\alpha}_{t}$ to the agents who then compute their $\boldsymbol{\beta}_{n,t}$ using $\boldsymbol{\beta}_{n,t}=\boldsymbol{A}^{\intercal}_{n}\boldsymbol{\alpha}$. Here $\boldsymbol{A}^{\intercal}_{n}$ denotes the $d$ rows corresponding to player $n$ in the matrix $\boldsymbol{A}^{\intercal}$. Our algorithm requires each player to know only its corresponding $d$ rows in $\boldsymbol{A}^{\intercal}$. Note that $\boldsymbol{\beta}=\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$. Depending on the application, this row can represent local parameters known to the player (e.g., the size of the jobs the player sends to the server), or that the manager sends these parameters to players at the beginning of the algorithm.
The main challenge in guaranteeing that the dynamics converge to a point where the linear constraints are satisfied is that the manager does not know the game, consisting of $\left\{r_{n}\right\}$ and $\left\{\mathcal{X}_{n}\right\}$. To overcome this, we propose an online approach where the manager learns how to adjust $\boldsymbol{\alpha}_{t}$ to satisfy the linear constraints by using the iteration in (8) with the control step-size sequence $\left\{\varepsilon_{t}\right\}$. We only assume that the manager can observe the instantaneous constraint violation $\boldsymbol{A}\boldsymbol{x}_{t-1}-\boldsymbol{\ell^{*}}$ at the start of turn $t$. This feedback maintains basic privacy for the players since the manager does not know their reward functions and does not even observe their individual actions.
Our algorithm is a two time-scale stochastic approximation algorithm. In the faster time-scale, dictated by the stepsize $\{\eta_{t}\}$, the agents update their action using stochastic gradient descent, taking a step towards the NE corresponding to $\boldsymbol{\alpha}_{t}$. In the slower timescale, dictated by the stepsize $\{\epsilon_{t}\}$, the manager observes the constraint violation based on the players’ current action updates $\boldsymbol{\alpha}_{t}$ to steer the player towards a desirable point.
Essentially, our algorithm is an iterative approach to solving the constrained optimization problem using Lagrange multipliers. We wish to solve the optimization problem with gradients given by $F(\boldsymbol{x})$ under the constraints $\boldsymbol{x}\in\mathcal{X}$ and $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$. The agents cannot ensure that $\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}}$ is satisfied for all $t$ without communicating, and hence the manager uses an iterative approach instead. Similar approaches have been used in [57, 58] for different problems.
Our players run stochastic gradient descent to optimize their rewards, and therefore are myopic and do not look to manipulate the manager using dynamic strategies. This behavior can be encountered with selfish players in a large-scale network or with cooperative players.
In a selfish setting, players are likely to maximize their rewards using a widespread online learning algorithm such as gradient descent [59, 10]. In particular, when the players do not know their reward functions, the small steps taken by SGD allow for gradual learning, as opposed to best-response dynamics that can have large jumps. Nevertheless, our techniques can be applied to other distributed algorithms that are guaranteed to converge to NE.
In a cooperative setting, the behavior of the players is not something to model but to design [60]. The algorithm is used as a distributed protocol programmed into the devices in the network (e.g., WiFi). The challenge is then not to control the selfish behavior of the players, but to guide them towards a more efficient NE they do not have enough information to find otherwise (i.e., players do not know which $\boldsymbol{\alpha}$ would satisfy the linear constraints at NE). For this purpose, gradient play is an efficient distributed algorithm, that can lead to an efficient NE when combined with the manager’s algorithm in (8).
The manager’s algorithm in (8) is robust against strategic manipulations of a small group of players since each player has a negligible effect on $\boldsymbol{A}\boldsymbol{x}_{t-1}$, which is its only way to impact (8). We conjecture that even a large group of players cannot manipulate (8) to increase their accumulated reward over time since any manipulated change to $\boldsymbol{\alpha}_{t}$ would be temporary and incurs losses to the players. Moreover, truthful reporting is not an issue since the players do not report anything, and the manager cannot observe their actions. Instead, the manager only needs to measure the instantaneous constraint violation.
3.1 Performance Guarantees
The manager wants to converge to a suitable $\boldsymbol{\alpha}$ such that the linear constraints in (6) are satisfied at the NE corresponding to $\boldsymbol{\alpha}$, i.e., converge to the set
$\mathcal{N}_{opt}=\{\boldsymbol{\alpha}\mid\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}\},$
where $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ denotes the NE corresponding to $\boldsymbol{\alpha}$.
We make the following assumption on our game to ensure that $\mathcal{N}_{opt}$ is non-empty, which is just Slater’s condition [61] for convex optimization.
Assumption 4.
The action set is of the form $\mathcal{X}=\{\boldsymbol{x}\mid h_{i}(\boldsymbol{x})\leq 0,i=1,\ldots,M\}$ for some $M$ where each $h_{i}(\boldsymbol{x})$ is continuously differentiable and convex and there exists $\boldsymbol{x}\in\text{int}(\mathcal{X})$ (i.e., $h_{i}(\boldsymbol{x})<0$, $i=1,\ldots,M$) such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$.
We also make the following standard assumptions on our step-size sequences [11].
Assumption 5.
The stepsizes $\{\eta_{t}\}$ and $\{\epsilon_{t}\}$ are of the form $\eta_{t}=1/(t+T_{1})^{\eta}$ and $\epsilon_{t}=1/(t+T_{2})^{\epsilon}$ for some $T_{1},T_{2}>0$ which satisfy:
  1. (a)
    $\{\eta_{t}\},\{\epsilon_{t}\}$ are monotonically decreasing sequences that are non-summable, but square summable.
  2. (b)
    $\lim_{t\to\infty}\frac{\epsilon_{t}^{2}}{\eta_{t}^{3}}=0$.
Our first result shows that $\boldsymbol{\alpha}_{t}$ converges to $\mathcal{N}_{opt}$ and hence $\boldsymbol{A}\boldsymbol{x}_{t}$ converges to $\boldsymbol{\ell^{*}}$.
Theorem 1.
Under Assumptions 1-5, $\boldsymbol{\alpha}_{t}$ converges to the set $\mathcal{N}_{opt}$, $\boldsymbol{x}_{t}$ converges to $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, and
$\lim_{t\rightarrow\infty}\boldsymbol{A}\boldsymbol{x}_{t}=\boldsymbol{\ell^{*}},$
with probability 1.
The following theorem gives finite-time convergence guarantees for our algorithm.
Theorem 2.
Under Assumptions 1-5,
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right).$
Therefore the best rate is obtained for $\eta=0.5+\delta/3$ and $\epsilon=0.75+\delta$:
$\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq O\left(t^{-0.25+\delta}\right),$
where $\delta$ is arbitrarily small.
The above theorem holds for all $t$ if we choose $T_{1}$ and $T_{2}$ to be sufficiently large and then $\mathbb{E}\left[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}\right]\leq C\left(\eta_{t}+\frac{1}{t\epsilon_{t}}\right)$ for all $t$, for some $C>0$.
A smaller value of $\epsilon$ gives a better rate. The optimal rate is then dictated by the assumptions that $\eta_{t}$ is square-summable ($\eta>0.5$) and $\lim_{t\uparrow\infty}\epsilon_{t}^{2}/\eta_{t}^{3}=0$ ($2\epsilon>3\eta$).
4 Convergence Analysis
In this section, we detail our proof strategy for Theorem 1 and 2. Throughout the proofs, we denote the gradient vector given the controlled input $\boldsymbol{\alpha}$ by $F(\boldsymbol{x},\boldsymbol{\alpha})=F(\boldsymbol{x})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$.
Our first lemma shows that the desired set $\mathcal{N}_{opt}$ is non-empty. The proof follows from the strong monotonicity of the game and Slater’s conditions (Assumption 4).
Lemma 3.
Under Assumptions 1 and 4, there exists $\boldsymbol{\alpha}$ such that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$, i.e., the set $\mathcal{N}_{opt}$ is non-empty.
Proof.
Let $\mathcal{G}(\boldsymbol{\alpha})$ be our $N$-player game with utility functions $u_{n}(\boldsymbol{x},\boldsymbol{\beta})$ and action set $\mathcal{X}$ where $\boldsymbol{\beta}=\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}$. Consider a modified $N$-player game $\mathcal{G}^{\prime}$ with utility function $u^{\prime}_{n}(\boldsymbol{x})=r_{n}(\boldsymbol{x})$ for $n\in\{1,\ldots,N\}$ and action set $\{\boldsymbol{x}\mid\boldsymbol{x}\in\mathcal{X},\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}\}$. Under the assumption that there exists $\boldsymbol{x}$ such that $\boldsymbol{A}\boldsymbol{x}=\boldsymbol{\ell^{*}}$ and $h_{i}(\boldsymbol{x})<0$ for all $i=1,\ldots,M$, Theorem 1 along with equations (3.3) and (3.6) from [8] states that there exists a NE $\hat{\boldsymbol{x}}\in\mathcal{X}$ of $\mathcal{G}^{\prime}$, and $\boldsymbol{\mu}\in\mathbb{R}^{M},\boldsymbol{\lambda}\in\mathbb{R}^{K}$ such that
$F(\hat{\boldsymbol{x}})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}+\boldsymbol{A}^{\intercal}\boldsymbol{\lambda}=0$
and
$\boldsymbol{A}\hat{\boldsymbol{x}}=\boldsymbol{\ell^{*}}\;\textrm{and}\;h_{i}(\hat{\boldsymbol{x}})\leq 0,\forall i\in\{1,2,\ldots,M\}.$
Let $\boldsymbol{\alpha}=-\boldsymbol{\lambda}$. Then the above equations can be written as
$F(\hat{\boldsymbol{x}},\boldsymbol{\alpha})+\sum_{i=1}^{M}\nabla h(\hat{\boldsymbol{x}})\mu_{i}=0$
and
$h_{i}(\hat{\boldsymbol{x}})\leq 0,\forall i\in\{1,2,\ldots,M\},$
which implies that $\hat{\boldsymbol{x}}$ is a NE for the game $\mathcal{G}(\boldsymbol{\alpha})$, and hence $\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\hat{\boldsymbol{x}}$. We also know that $\boldsymbol{A}\hat{\boldsymbol{x}}=\boldsymbol{\ell^{*}}$. Hence there exists $\boldsymbol{\alpha}$ such that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$, showing that $\mathcal{N}_{opt}$ is non-empty.
One of the main technical challenges in our analysis stems from the constrained actions represented by the convex and compact action sets. Indeed, with no constraints on the actions, and under certain assumptions on the matrix $\boldsymbol{A}$, we can show that $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is strongly monotone instead of co-coercive (Lemma 4 (b)). This in turn shows that the mapping $g(\boldsymbol{\alpha})$ in Definition 2 is contractive. In this simplified case, both time-scales are fixed point iterations with contractive mappings and results from [11] can be applied. Therefore, the action sets call for a novel two time scale convergence analysis.
4.1 Non-expansive Mapping
In this subsection, we prove the non-expansiveness of the following mapping that is central to our analysis.
Definition 2.
Define the map $g(\cdot):\mathbb{R}^{K}\mapsto\mathbb{R}^{K}$ as
$g(\boldsymbol{\alpha})=\boldsymbol{\alpha}+\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}}),$
where $\gamma=\min\{\mu/\|\boldsymbol{A}\|^{2},1\}$ is a constant chosen to satisfy $02\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2}>$ 0. Note that a fixed point of $g(\boldsymbol{\alpha})$ will satisfy our objective of $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})=\boldsymbol{\ell^{*}}$.
To establish the non-expansiveness of $g$, we first need to show two important properties of the NE corresponding to each $\boldsymbol{\alpha}\in\mathbb{R}^{K}$. The first property is that $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is Lipschitz continuous in $\boldsymbol{\alpha}$ and the second property is that the weighted load vector $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is co-coercive in $\boldsymbol{\alpha}$. The proofs for these two results follow from the strongly monotone nature of the game by using properties of variational inequalities (VI).
Lemma 4.
Let $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ be the unique NE given $\boldsymbol{\alpha}\in\mathbb{R}^{K}$. Then,
  1. (a)
    There exists a constant $L_{0}$ such that for all $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$,
    $\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|\leq L_{0}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|.$ (9)
  2. (b)
    For all $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$,
    $\langle\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\rangle$
    $\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\leq-\mu\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2},$ (10)
    where $\mu$ is the monotonicity parameter of the game.
Proof.
For a fixed $\boldsymbol{\alpha}$, $-F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone on $\mathcal{X}$ and is Lipschitz continuous by Assumptions 1 and 2 respectively. Then [51, Theorem 2.3.3] states that for all $\boldsymbol{x}\in\mathcal{X}$, for some $L^{\prime}$
$\|\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha})\|\leq L^{\prime}\|\boldsymbol{x}-\Pi_{\mathcal{X}}(\boldsymbol{x}+F(\boldsymbol{x},\boldsymbol{\alpha}))\|.$
Hence for $\boldsymbol{\alpha}_{2},\boldsymbol{\alpha}_{1}\in\mathbb{R}^{K}$, we have
$\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|$
$\leq L^{\prime}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\|$
=$\stackrel{{\scriptstyle(a)}}{{=}}L^{\prime}\Big{\|}\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}))$
$\;\;\;\;\;\;\;\;\;-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}))\Big{\|}$
$\leq L^{\prime}\|F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2})-F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1})\|$
$=L^{\prime}\|\boldsymbol{A}^{\intercal}(\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1})\|$
$\leq L_{0}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|,$
where $L_{0}=L^{\prime}\|\boldsymbol{A}^{\intercal}\|$. Here equality (a) follows from [51, Proposition 1.5.8] which says
$\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})=\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})+F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2})).$
This completes the proof for part (a) of Lemma 4.
Now let $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$ Since $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})$ is the NE corresponding to $\boldsymbol{\alpha}_{1}$, using [51, Proposition 1.4.2], we have for every $\boldsymbol{x}\in\mathcal{X}$ that $\left\langle\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\leq 0$. For $\boldsymbol{x}=\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})$
$\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\leq 0.$
Similarly, we also have
$\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}\right)\right\rangle\leq 0.$
By combining the above two inequalities we obtain
$\left\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}\right)-F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\right\rangle\geq 0.$ (11)
Then
$-\mu\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$\stackrel{{\scriptstyle(a)}}{{\geq}}\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{1}\right)-F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\rangle$
$=\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2}),\boldsymbol{\alpha}_{2}\right)-F\left(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{1}\right)\rangle$
$\;\;\;+\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{2}-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{1}\rangle$
$\stackrel{{\scriptstyle(b)}}{{\geq}}\langle\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{2}-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{1}\rangle$
$=(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}))^{\intercal}\boldsymbol{A}^{\intercal}(\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1})$
$=(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}))^{\intercal}(\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1})$
$=\left\langle\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\right\rangle.$
Here inequality (a) uses that $F(\boldsymbol{x},\boldsymbol{\alpha})$ is strongly monotone in $\boldsymbol{x}$ with parameter $\mu$ and (b) uses (11). ∎
The next lemma shows that the map $g(\cdot)$ is non-expansive.
Lemma 5.
For any $\boldsymbol{\alpha}_{1},\boldsymbol{\alpha}_{2}\in\mathbb{R}^{K}$, the function $g(\cdot)$ satisfies,
$\|g(\boldsymbol{\alpha}_{2})-g(\boldsymbol{\alpha}_{1})\|\leq\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|,$ (12)
i.e., the map $g(\cdot)$ is non-expansive.
Proof.
For any $\boldsymbol{\alpha}_{2},\boldsymbol{\alpha}_{1}\in\mathbb{R}^{K}$,
$\|g(\boldsymbol{\alpha}_{2})-g(\boldsymbol{\alpha}_{1})\|^{2}$
$=\|\boldsymbol{\alpha}_{2}+\gamma\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{\alpha}_{1}-\gamma\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$=\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}+\gamma^{2}\left\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\right\|^{2}$
$\;\;\;+2\gamma\left\langle\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1}),\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\right\rangle$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}+\gamma^{2}\|\boldsymbol{A}\|^{2}\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$\;\;\;-2\mu\gamma\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$\leq\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}-(2\mu\gamma-\gamma^{2}\|\boldsymbol{A}\|^{2})\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{2})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{1})\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\|\boldsymbol{\alpha}_{2}-\boldsymbol{\alpha}_{1}\|^{2}.$
Here inequality (a) follows from part (b) of Lemma 4 and inequality (b) follows from the assumption on $\gamma$ that $2\mu\gamma>\gamma^{2}\|\boldsymbol{A}\|^{2}$. Hence $g(\cdot)$ is non-expansive. ∎
The next lemma shows that the iterate $\boldsymbol{\alpha}_{t}$ has bounded expectation for all $t$. This lemma is required both for proving the asymptotic convergence result in Theorem 1, which requires almost-sure boundedness of iterates and to obtain a better convergence rate in Theorem 2.
Lemma 6.
Under Assumptions 1-5,
$\mathbb{E}[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}]\leq C_{0},$
for some $C_{0}>0$, $\boldsymbol{\alpha^{*}}\in\mathcal{N}_{opt}$ and all $t>0$.
Proof.
Let $\boldsymbol{\alpha^{*}}\in\mathcal{N}_{opt}$. We first give a recursive bound on the term $\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$. For this, note that
$\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
$=\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1}))-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\|\Pi_{\mathcal{X}}(\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1}))$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;-\Pi_{\mathcal{X}}(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1}))\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\|\boldsymbol{x}_{t}+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
$=\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}(F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{M}_{t+1})$
$\;\;+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
$=\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}$ (13a)
$\;\;+\|\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$ (13b)
$\;\;+2\big{\langle}\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t}),$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\big{\rangle}.$ (13c)
Here equality (a) uses that $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\in\mathcal{X}$ and inequality (b) follows from since the projection operator $\Pi_{\mathcal{X}}(\cdot)$ for any convex set $\mathcal{X}$ is non-expansive.
We first simplify the term (13a):
$\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}$
$=\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\eta_{t}^{2}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}$
$\;\;+2\eta_{t}\langle\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\rangle.$
The last term is bounded as follows.
$\langle\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\rangle$
$\leq\langle\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})-F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t})\rangle$
$\leq-\mu\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}.$
Here the first inequality follows since $\langle\boldsymbol{x}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}),F(\boldsymbol{x^{*}}(\boldsymbol{\alpha}),\boldsymbol{\alpha})\rangle\leq 0$, for all $\boldsymbol{\alpha}\in\mathbb{R}^{K}$ and $\boldsymbol{x}\in\mathcal{X}$, since $\boldsymbol{x^{*}}(\boldsymbol{\alpha})$ is a NE (Proposition 1.4.2 in [51]) and the second inequality follows from the strong monotonicity property of the game. We also have that
$\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}=\|F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}\|^{2}$
$=\|F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha^{*}}-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha}_{t}+\boldsymbol{A}^{\intercal}\boldsymbol{\alpha^{*}}\|^{2}$
$\leq 2\|F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha^{*}}\|^{2}+2\|\boldsymbol{A}^{\intercal}\|^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}$
$\leq c_{1}^{2}+c_{2}^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}.$
where the last inequality follows for some $c_{1}$ and $c_{2}$ as $\|F(\boldsymbol{x}_{t})-\boldsymbol{A}^{\intercal}\boldsymbol{\alpha^{*}}\|$ is bounded by some constant as $\boldsymbol{x}_{t}\in\mathcal{X}$, where $\mathcal{X}$ is compact. Coming back to (13a), we have
$\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}$
$\leq(1-2\mu\eta_{t})\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\eta_{t}^{2}c_{1}^{2}+\eta_{t}^{2}c_{2}^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}.$
For (13b), note that
$\mathbb{E}[\|\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}\mid\mathcal{F}_{t}]$
$=\eta_{t}^{2}\mathbb{E}[\|\boldsymbol{M}_{t+1}\|^{2}\mid\mathcal{F}_{t}]+\mathbb{E}[\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}\mid\mathcal{F}_{t}]$
$\;\;+2\eta_{t}\mathbb{E}[\langle\boldsymbol{M}_{t+1},\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\rangle\mid\mathcal{F}_{t}]$
=$\stackrel{{\scriptstyle(a)}}{{=}}\eta_{t}^{2}\mathbb{E}[\|\boldsymbol{M}_{t+1}\|^{2}\mid\mathcal{F}_{t}]+\|\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}$
$\stackrel{{\scriptstyle(b)}}{{\leq}}\eta_{t}^{2}M_{c}+L_{0}^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha}_{t+1}\|^{2}=\eta_{t}^{2}M_{c}+L_{0}^{2}\epsilon_{t}^{2}\left\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\right\|^{2}$
$\stackrel{{\scriptstyle(c)}}{{\leq}}\eta_{t}^{2}M_{c}+\epsilon_{t}^{2}c_{3}^{2}.$
Here equality (a) follows from the fact that $\alpha_{t+1}$ is $\mathcal{F}_{t}$ measurable and $\boldsymbol{M}_{t+1}$ is a Martingale difference sequence, hence making the term $\mathbb{E}[\langle\boldsymbol{M}_{t+1},\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\rangle\mid\mathcal{F}_{t}]=0$. Inequality (b) follows from Lemma 4 part (a), and (c) follows from $\boldsymbol{x}_{t}\in\mathcal{X}$ for some constant $c_{3}$.
Finally, for the third term (13c):
$2\mathbb{E}\big{[}\big{\langle}\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t}),$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\big{\rangle}\mid\mathcal{F}_{t}\big{]}$
$=2\big{\langle}\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t}),\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\big{\rangle}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|L_{0}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha}_{t+1}\|$
$\stackrel{{\scriptstyle}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|\epsilon_{t}L_{0}\left\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\right\|$
$\;\;+2\eta_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|\epsilon_{t}L_{0}\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|$
$\stackrel{{\scriptstyle(b)}}{{\leq}}2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|\epsilon_{t}c_{3}+2\eta_{t}\epsilon_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|c_{3}.$
Here the first equality uses that $\boldsymbol{M}_{t+1}$ is a martingale difference sequence and $\boldsymbol{x}_{t}$, $\boldsymbol{\alpha}_{t}$ and $\boldsymbol{\alpha}_{t+1}$ are $\mathcal{F}_{t}$ measurable. Inequality (a) follows from the Cauchy-Schwarz inequality and the Lipschitz continuity of the map $\boldsymbol{x^{*}}(\cdot)$ (Lemma 4) and (b) uses boundedness of $\boldsymbol{x}_{t}$ and similar to previous expansions.
To bound the term $2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|\epsilon_{t}c_{3}$, we use $(2ab\leq a^{2}+b^{2})$ where $a=\sqrt{\mu\eta_{t}\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}/2}$ and $b=\sqrt{(\epsilon_{t}^{2}/\eta_{t})(2c_{3}^{2}/\mu)}$. To bound the term $2\eta_{t}\epsilon_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|c_{3}$, we use the same inequality with $a=\eta_{t}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|$ and $b=\epsilon_{t}c_{3}$. This yields
$2\mathbb{E}\big{[}\big{\langle}\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})+\eta_{t}F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t}),$
$\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\eta_{t}\boldsymbol{M}_{t+1}+\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\big{\rangle}\mid\mathcal{F}_{t}\big{]}$
$\leq\frac{\mu}{2}\eta_{t}\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\frac{\epsilon_{t}^{2}}{\eta_{t}}2c_{3}^{2}/\mu+\eta_{t}^{2}\|F(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\|^{2}+\epsilon_{t}^{2}c_{3}^{2}$
$\leq\frac{\mu}{2}\eta_{t}\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\frac{\epsilon_{t}^{2}}{\eta_{t}}2c_{3}^{2}/\mu$
$\;\;\;+\eta_{t}^{2}c_{1}^{2}+\eta_{t}^{2}c_{2}^{2}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}+\epsilon_{t}^{2}c_{3}^{2}$
We have now bounded all three terms from (13). Combining them after taking expectation gives us:
$\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$
$\leq(1-\frac{3}{2}\mu\eta_{t})\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]+\eta_{t}^{2}(2c_{1}^{2}+M_{c})$
$\;\;\;+2\epsilon_{t}^{2}c_{3}^{2}+\frac{\epsilon_{t}^{2}}{\eta_{t}}2c_{3}^{2}/\mu+2\eta_{t}^{2}c_{2}^{2}\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right].$
Our assumption that $\lim_{t\to\infty}\epsilon_{t}^{2}/\eta_{t}^{3}\rightarrow 0$ implies that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ for some $C_{1}>0$ and $t>T_{0}$. Then for all $t>T_{0}$:
$\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$ $\leq(1-1.5\mu\eta_{t})\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$
$+$ $c_{4}\eta_{t}^{2}+2\eta_{t}^{2}c_{2}^{2}\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right],$ (14)
for $c_{4}=2c_{1}^{2}+M_{c}+2C_{1}c_{3}^{2}+2C_{1}c_{3}^{2}/\mu$.
We next give a recursive bound on $\mathbb{E}[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}]$. Let $\epsilon_{t}^{\prime}=\epsilon_{t}/\gamma$. Then iteration (8) can be written as
$\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}+\epsilon_{t}^{\prime}(g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t})+\epsilon_{t}^{\prime}\boldsymbol{A}\gamma(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})),$
where $g(\boldsymbol{\alpha})=\boldsymbol{\alpha}+\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}})$. For ease of notation define the sequences $p_{t}=g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}$ and $q_{t}=\gamma\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))$. Let $\boldsymbol{\alpha^{*}}\in\mathcal{N}_{opt}$ be a fixed point of $g(\cdot)$. Then
$\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}$
$=\|\boldsymbol{\alpha}_{t}+\epsilon_{t}^{\prime}(g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}+q_{t})-\boldsymbol{\alpha^{*}}\|^{2}$
$=\|(1-\epsilon_{t}^{\prime})(\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}})+\epsilon^{\prime}_{t}(g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}+q_{t})\|^{2}.$
Using Corollary 2.14 from [61], we obtain
$\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}$ $\leq(1-\epsilon_{t}^{\prime})\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}$
$\;\;\;+\epsilon_{t}^{\prime}\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}+q_{t}\|^{2}$
$\;\;\;-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\|p_{t}+q_{t}\|^{2}.$ (15)
We bound the second term as follows:
$\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}+q_{t}\|^{2}$
$=\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}\|^{2}+\|q_{t}\|^{2}+2\langle g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}},q_{t}\rangle$
$\leq\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}+\|q_{t}\|^{2}+2\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha^{*}}\|\|q_{t}\|$
$\leq\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}+\|q_{t}\|^{2}+2\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|\|q_{t}\|$
Here the first inequality stems from the non-expansive nature of map $g(\cdot)$ and the final inequality uses the fact that $g(\boldsymbol{\alpha^{*}})=\boldsymbol{\alpha^{*}}$ along with the non-expansive nature of $g(\cdot)$. Now
$2\epsilon_{t}^{\prime}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|\|q_{t}\|\leq 2\epsilon_{t}\|\boldsymbol{A}\|\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|$
$\leq\mu\eta_{t}/2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\frac{2\|\boldsymbol{A}\|^{2}\epsilon_{t}^{2}}{\mu\eta_{t}}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}.$
Then returning to (4.1),
$\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\leq\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}+\epsilon_{t}^{\prime}\gamma^{2}\|\boldsymbol{A}\|^{2}\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}$
$\;\;\;+\mu\eta_{t}/2\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\frac{2\|\boldsymbol{A}\|^{2}\epsilon_{t}^{2}}{\mu\eta_{t}}\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}$
$\;\;\;-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\|p_{t}+q_{t}\|^{2}.$
Again using our assumption that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ for some appropriate $C_{1}>0$ and $t>T_{0}$, and taking expectation we have
$\mathbb{E}\left[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
$\leq\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+\mu\eta_{t}\mathbb{E}\left[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}\right]$
$\;\;\;+2\eta_{t}^{2}c_{2}^{2}\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\mathbb{E}[\|p_{t}+q_{t}\|^{2}].$ (16)
Adding (4.1) and (4.1), we get
$\mathbb{E}\left[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
$\leq(1+4c_{2}^{2}\eta_{t}^{2})\mathbb{E}\left[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+c_{4}\eta_{t}^{2}.$
Expanding the recursion for $t>T_{0}$ gives us
$\mathbb{E}\left[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
$\leq\mathbb{E}\left[\|\boldsymbol{x}_{T_{0}}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{T_{0}})\|^{2}+\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}\right]\prod_{i=T_{0}}^{t-1}(1+4c_{2}^{2}\eta_{i}^{2})$
$\;\;\;+c_{4}\sum_{i=T_{0}}^{t-1}\eta_{i}^{2}\prod_{j=i+1}^{t-1}(1+4c_{2}^{2}\eta_{j}^{2})$
$\leq\mathbb{E}\left[\|\boldsymbol{x}_{T_{0}}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{T_{0}})\|^{2}+\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}\right]e^{\left(4c_{2}^{2}\sum_{i=T_{0}}^{t-1}\eta_{i}^{2}\right)}$
$\;\;\;+c_{4}\sum_{i=T_{0}}^{t-1}\eta_{i}^{2}e^{\left(4c_{2}^{2}\sum_{j=i+1}^{t-1}\eta_{j}^{2}\right)}.$
Here the final inequality uses that $(1+x)\leq e^{x}$ for any $x\in\mathbb{R}$. Now we use our assumption that $\eta_{t}$ is square-summable. Hence, $\exp(4c_{2}^{2}\sum_{i=T_{0}}^{t-1}\eta_{i}^{2})$ and $\sum_{i=T_{0}}^{t-1}\eta_{i}^{2}\exp\left(4c_{2}^{2}\sum_{j=i+1}^{t-1}\eta_{j}^{2}\right)$ are bounded by constants. Finally, $T_{0}$ is a constant dependent only on the stepsize sequence, and hence $\|\boldsymbol{x}_{T_{0}}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{T_{0}})\|^{2}+\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}$ can be bounded by a constant using the discrete Gronwall’s inequality [62]. This finally gives us that for some constant $C_{0}$
$\mathbb{E}\left[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]\leq C_{0},$
which completes the proof for Lemma 6. ∎
The above three lemmas are sufficient for proving Theorem 1 (see Appendix). We use the properties of two time-scale stochastic approximation to show that $(\boldsymbol{x}_{t},\boldsymbol{\alpha}_{t})\to(\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t})$. This allows us to analyze the slower iteration, i.e., the control updates (8) in the limit $\boldsymbol{x}_{t}\to\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$. We rewrite the iteration as a fixed point iteration with the mapping $g(\cdot)$. Since $g(\cdot)$ is non-expansive, this iteration is a version of the inexact KM algorithm [61]. We study its ODE limit and show convergence to the set of fixed points.
4.2 Finite Time Convergence Guarantees
In this subsection we detail how to obtain our finite time convergence guarantees. This result also uses the non-expansiveness on $g$, but requires a more nuanced argument that bounds the error as follows
$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x}_{t}-\boldsymbol{\ell^{*}}\|^{2}]$ $\leq 2\mathbb{E}[\|\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))\|^{2}]$
$\;\;\;\;\;\;\;+2\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}].$
We first bound the mean square distance between $\boldsymbol{x}_{t}$ and $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, i.e., the NE corresponding to $\boldsymbol{\alpha}_{t}$. The key to proving this lemma is the strong monotonicity of the game which gives a recursive bound on $\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$. This recursive bound gives a rate of $O(\eta_{t})$ when solved.
Lemma 7.
Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime}>0$ and $T_{0}$ such that for all $t>T_{0}$:
$\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]\leq C^{\prime}\eta_{t}.$
Proof.
Returning to the the recursive bound (4.1) from the proof for Lemma 6, and using Lemma 6 that shows that $\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]\leq C_{0}$, we now obtain the recursion
$\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$ $\leq(1-1.5\mu\eta_{t})\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$
$+\eta_{t}^{2}(c_{4}+2c_{2}^{2}C_{0}).$
For $\eta_{t}=(t+T_{1})^{-\eta}$ with $10<\eta<$ 1, this recursion yields [12]
$\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]\leq C^{\prime}\eta_{t},$
for $t>T_{0}$ for some constant $C^{\prime}$. ∎
The next lemma bounds the mean squared distance between $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and $\boldsymbol{\ell^{*}}$. Note that we need to bound $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ as $g(\boldsymbol{\alpha})-\boldsymbol{\alpha}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}})$. We prove the bound in the following lemma using Lemma 7 and a technique similar to the bound for (inexact) KM algorithm. If we could assume that $\epsilon_{t}>\epsilon^{*}$ for some $\epsilon^{*}$, then $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ is known to decay at the rate $o(1/t)$ [63]. Since this assumption is not true in our case, we use the assumptions on the stepsize sequences to get a convergence rate of $O(1/(t\epsilon_{t}))$.
Lemma 8.
Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime\prime}>0$ and $T_{0}>0$ such that for all $t>T_{0}$.
$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]\leq\frac{C^{\prime\prime}}{t\epsilon_{t}}.$
Proof.
Define $p_{t}=g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}$, $q_{t}=\gamma\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))$ and $\epsilon_{t}^{\prime}=\epsilon_{t}/\gamma$. Returning to the recursive bound (4.1) from the proof of Lemma 6, and now using Lemma 6 and Lemma 7,
$\mathbb{E}\left[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
$\leq\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+\eta_{t}^{2}c_{5}^{2}-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\mathbb{E}[\|p_{t}+q_{t}\|^{2}]$ (17)
for some constant $c_{5}$. For the third term above, we note that
$-\|p_{t}+q_{t}\|^{2}$ $=-\|p_{t}\|^{2}-\|q_{t}\|^{2}-2\langle p_{t},q_{t}\rangle$
$\leq-\|p_{t}\|^{2}-\|q_{t}\|^{2}+2\|p_{t}\|\|q_{t}\|$
$\leq-\|p_{t}\|^{2}+c_{6}\|q_{t}\|.$
Note that $p_{t}=g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}})$. Hence the last inequality follows from the compactness of $\mathcal{X}$ for some constant $c_{6}>0$. This implies $\mathbb{E}\left[-\|p_{t}+q_{t}\|^{2}\right]\leq\mathbb{E}\left[-\|p_{t}\|^{2}\right]+2c_{7}\sqrt{\eta_{t}},$ for some constant $c_{7}>0$. This inequality follows from Lemma 7 and since $\mathbb{E}[\|q_{t}\|]^{2}\leq\mathbb{E}[\|q_{t}\|^{2}]$.
Using our assumption that $\epsilon_{t}$ goes to zero, we have that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$. For the rest of the proof, we pick $T_{0}$ to be large enough such that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$, $\epsilon_{t}<\gamma/2$ and $\eta_{t}<C_{2}$ for all $t>T_{0}$, for appropriate $C_{1}$ and $C_{2}$. Returning to (4.2), for all $t>T_{0}$, we have
$\mathbb{E}\left[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
$\leq\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+\eta_{t}^{2}c_{5}^{2}+\epsilon_{t}^{\prime}\sqrt{\eta_{t}}c_{7}-(\epsilon_{t}^{\prime}/2)\mathbb{E}[\|p_{t}\|^{2}].$
This gives us
$(\epsilon_{t}^{\prime}/2)\mathbb{E}[\|p_{t}\|^{2}]$ $\leq\mathbb{E}[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}]-\mathbb{E}[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}]$
$\;\;\;+\eta_{t}^{2}c_{5}^{2}+\epsilon_{t}^{\prime}\sqrt{\eta_{t}}c_{7}.$
Summing from $i=T_{0}$ to $t$ gives us
$\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\mathbb{E}[\|p_{i}\|^{2}]$ $\leq\mathbb{E}[\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}]-\mathbb{E}[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}]$
$\;\;+\sum_{i=T_{0}}^{t}(\eta_{i}^{2}c_{5}^{2}+\epsilon_{i}^{\prime}\sqrt{\eta_{i}}c_{7}).$ (18)
The left hand term in the above inequality depends on $\|p_{i}\|^{2}$ for all $i\in\{T_{0},\ldots,t\}$. We next show a bound for $\|p_{i}\|^{2}$ in terms of $\|p_{t}\|^{2}$ for all $i\in\{T_{0},\ldots,t\}$. For this, we first recall that $\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}+\epsilon_{t}^{\prime}(p_{t}+q_{t}).$ Next,
$\|p_{t+1}\|^{2}=\|p_{t}\|^{2}+\|p_{t+1}-p_{t}\|^{2}+2\langle p_{t+1}-p_{t},p_{t}\rangle$
$\leq\|p_{t}\|^{2}+\|p_{t+1}-p_{t}\|^{2}+\frac{2}{\epsilon_{t}^{\prime}}\langle p_{t+1}-p_{t},\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}-\epsilon_{t}^{\prime}q_{t}\rangle$
Now,
$2\langle p_{t+1}-p_{t},\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\rangle$
$=2\langle g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t+1}+\boldsymbol{\alpha}_{t},\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\rangle$
$=2\langle g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\rangle-2\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}$
$=\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}-2\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}$
$\;\;-\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t+1}+\boldsymbol{\alpha}_{t}\|^{2}$
$=\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})\|^{2}-\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}-\|p_{t+1}-p_{t}\|^{2}$
$\leq-\|p_{t+1}-p_{t}\|^{2}.$
Here the final inequality follows from the non-expansive nature of $g(\cdot)$. This gives us
$\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}-\frac{1-\epsilon_{t}^{\prime}}{\epsilon_{t}^{\prime}}\|p_{t+1}-p_{t}\|^{2}-2\langle p_{t+1}-p_{t},q_{t}\rangle$
$=\|p_{t}\|^{2}-\frac{1-\epsilon_{t}^{\prime}}{\epsilon_{t}^{\prime}}\left\|p_{t+1}-p_{t}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}q_{t}\right\|^{2}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}\|q_{t}\|^{2}$
$\leq\|p_{t}\|^{2}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}\|q_{t}\|^{2}.$
Using the assumption that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$, this gives us $\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}+2\epsilon_{t}^{\prime}\|q_{t}\|^{2}.$ Taking expectation and using Lemma 7, this results in the bound
$\mathbb{E}[\|p_{i}\|^{2}]$ $\geq\mathbb{E}[\|p_{t}\|^{2}]-\sum_{j=i}^{t-1}2\epsilon_{j}^{\prime}\mathbb{E}\|q_{j}\|^{2}$
$\geq\mathbb{E}[\|p_{t}\|^{2}]-\sum_{j=i}^{t-1}2c_{8}\epsilon_{j}^{\prime}\eta_{j},$
for some constant $c_{8}>0$. Returning to (4.2), we now have
$\mathbb{E}[\|p_{t}\|^{2}]\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\leq\mathbb{E}[\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}]$
$\;\;+\sum_{i=T_{0}}^{t}(\eta_{i}^{2}c_{5}^{2}+\epsilon_{i}^{\prime}\sqrt{\eta_{i}}c_{7})+2c_{8}\sum_{i=T_{0}}^{t}\epsilon_{i}^{\prime}\sum_{j=i}^{t-1}\epsilon_{j}^{\prime}\eta_{j}.$ (19)
Now we use our assumption that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ for all $t>T_{0}$. Then $\epsilon_{i}^{\prime}\sqrt{\eta_{i}}=O(\eta_{i}^{2})$ and hence the sequence $\{\epsilon_{i}^{\prime}\sqrt{\eta_{i}}\}$ is summable. Finally for the double summation, we use the integral test and Assumption 5 to obtain
$\sum_{i=T_{0}}^{t}\epsilon_{i}^{\prime}\sum_{j=i}^{t-1}\epsilon_{j}^{\prime}\eta_{j}\leq\frac{1}{\gamma^{2}}\sum_{i=T_{0}}^{\infty}\frac{1}{(i+T_{2})^{\epsilon}}\sum_{j=i}^{\infty}\frac{1}{(j+T_{3})^{\epsilon+\eta}}$
$\leq\sum_{i=T_{0}}^{\infty}\frac{1}{(i+T_{2})^{\epsilon}}\frac{c_{9}}{(i+T_{3})^{\epsilon+\eta-1}}\leq\sum_{i=T_{0}}^{\infty}\frac{c_{10}}{(i+T_{3})^{4\eta-1}}.$
Here $c_{9}$ and $c_{10}$ are constants and $T_{3}=\max\{T_{1},T_{2}\}$. The final inequality follows from the fact that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ and hence $2\epsilon\geq 3\eta$. Since $\eta_{t}$ is square summable, $\eta>0.5$ and hence $14\eta-1>$ 1, so the summation is bounded by a constant. Hence all terms in the right-hand side of (4.2) are bounded by a constant. Then for some $c_{11}>0$, we have
$\mathbb{E}[\|p_{t}\|^{2}]\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\leq c_{11}.$
Now $\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\geq(t-T_{0}+1)\epsilon_{t}^{\prime}/2$. This finally gives us the bound $\mathbb{E}[\|p_{t}\|^{2}]\leq C^{\prime\prime}/(t\epsilon_{t}),$ for all $t>T_{0}$. ∎
5 Simulation Results
(a) Loads for all resources
(b) Loads for 3 resources
Figure 2: Weighted Load Balancing in Resource Allocation Games
(a) Quadratic Cost
(b) Sum of rewards of all players
Figure 3: Quadratic Global Objective
In this section, we simulate the two applications discussed in Section 2.1. In all experiments, we use the step-size sequence $\eta_{t}=\frac{1}{\left(t+1\right)^{0.5}}$ and the control step-size sequence $\varepsilon_{t}=\frac{1}{\left(t+1\right)^{0.75}}$. We add Gaussian noise with mean 00 and variance $\sigma^{2}=\frac{1}{4}$ to the gradients of each player. We ran 100 realizations for each experiment and plotted the average result along with the standard deviation region, which was always small. In each realization, $\alpha_{0}^{k}$ and $x_{n,0}^{k}$ were chosen uniformly and independently at random on $\left[0,2\right]$ and $\left[0,0.1\right]$, respectively.
We pick a resource allocation game (Figure 2) inspired by DSM in electricity grid [64]. We keep $N=1000$ and $d=24$. We choose the action set for each player $n$ as $\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}.$ Here $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$, and $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $5$ and variance $1$. We use the reward function $r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$ Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$. We also defined $s_{i}=\sum_{n=1}^{N}a_{n,i}^{(i)}x_{n}^{i}$, i.e., the weighted load at resource $i$. The coefficients $a_{n,i}^{(i)}$ are i.i.d. samples from a Gaussian distribution with mean $1$ and variance $0.2$. The objective is $s_{i}=\ell_{i}^{*}$ where $\ell_{i}^{*}$ are i.i.d. samples from a Gaussian distribution with mean $35$ and variance $5$.
In Fig. 2(a), we plot the weighted load for each resource along with their target loads. We can observe that the system converges to a NE where the loads are balanced. In Fig. 2(b), we plot the weighted load for the first 3 resources. To study the effectiveness of our algorithm, we also plot the loads in the case of a uncontrolled system., i.e., where $\boldsymbol{\alpha}=\boldsymbol{0}$. It can be observed that if the system is uncontrolled, some resources are under-utilized while others are over-utilized.
For the quadratic global objective (Figure 3), we consider a quadratic game. In a quadratic game, the reward of each player is of the form $r_{n}\left(\boldsymbol{x}\right)=\boldsymbol{x}^{T}\boldsymbol{Q_{n}}\boldsymbol{x}+\boldsymbol{c_{n}}\boldsymbol{x}_{n}$ for some negative definite matrix $\boldsymbol{Q_{n}}\in\mathbb{R}^{Nd\times Nd}$ and $\boldsymbol{c_{n}}\in\mathbb{R}^{d}$.
We work with $N=100$ players and $d=5$, and action sets $\{\boldsymbol{x}_{n}\mid\sum_{i=1}^{d}x_{n}^{i}\leq 10\}$ for each player $n$. Recall that the global cost is given by $\boldsymbol{x}^{\top}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$. To generate $\boldsymbol{G}$, we first generate random $\boldsymbol{H_{0}}\in\mathbb{R}^{Nd\times(Nd-5)}$ with each element of $\boldsymbol{H_{0}}$ sampled independently from a standard Gaussian distribution. We use $\boldsymbol{G}=\boldsymbol{H_{0}}\boldsymbol{H_{0}}^{T}$. A random $\boldsymbol{\rho}$ is obtained with each element sampled from a Gaussian distribution with mean 0 and variance $0.5$. The above matrices $\boldsymbol{Q_{n}}$ are obtained by sampling $\boldsymbol{H_{n}}\in\mathbb{R}^{Nd\times Nd}$ and using $\boldsymbol{Q_{n}}=\boldsymbol{H_{n}}\boldsymbol{H_{n}}^{T}$. The vectors $\boldsymbol{c_{n}}$ are sampled in a similar way as $\boldsymbol{\rho}$.
In Fig. 3(a), we observe that our algorithm is able to converge to a lower quadratic cost as compared to a system which is left uncontrolled. In Fig. 3(b), we plot the sum of rewards for our algorithm. This is compared with the sum of rewards if the global objective was directly optimized for, i.e., each player updates their action based on $\nabla\Phi(\boldsymbol{x})$ where $\Phi(\boldsymbol{x})$ is the quadratic cost. Our algorithm provides a much higher sum of rewards after convergence. The two plots together show that our algorithm can be used as a method to optimize for a global cost while still trying to maximize the rewards for players.
6 Conclusion
We studied how to control strongly monotone games where the game structure is unknown to the manager. The goal of the controller is to ensure the game satisfies target linear constraints at NE. Our algorithm only requires the manager to observe the instantaneous constraint violation. We prove that our simple online algorithm guarantees convergence to a point $\boldsymbol{x^{*}}$ which satisfies the linear constraints $\boldsymbol{A}\boldsymbol{x^{*}}=\boldsymbol{l}^{*}$ by adjusting control inputs $\boldsymbol{\alpha}$ in real-time. Further, we show that the convergence rate for our algorithm is $O(t^{-1/4})$. Our algorithm is a two time-scale algorithm, where the algorithm does not wait for the players to converge to a NE but uses the feedback from every turn. This is an encouraging result since our algorithm is easy to implement and maintains the users’ privacy. The manager does not observe individual actions, and players do not need to reveal their reward functions, which might even be unknown to them.
Technically, we analyze how the NE of the game changes when the manager changes $\boldsymbol{\alpha}_{t}$. Our algorithm essentially controls the unknown game such that its NE becomes desirable from an optimization point of view. Since convergence to NE is relatively fast, this has the potential to accelerate general cooperative multi-agent optimization [60].
Controlling unknown games with bandit feedback is a new online learning paradigm. The game to be controlled is a non-stationary but highly structured bandit. Viewing game control as a bandit problem can result in new exciting research questions, and new powerful game control techniques. Generalizing our game control setting beyond linear constraints can help tackle a wider class of problems.